the length of a rectangle is 3 times the width, and the perimeter is 22. Find the deminsions of the rectangle.
Answer by jojo14344(1512)

Given:
Width (W) = x (unknown)
Length (L) = 3 times the width = 3W = 3x
Perimeter (P) = 22
Formula:
P = 2(L+W)
Substitution,
22 = 2 ( 3W + W )
22 = 2 ( 3x + x)
22/2 = 4x
x= 2.75 ( width )
Therefor the length (L) = 3 W = 3 (2.75)
L = 8.25
To check:
P = 2 (L + W)
22= 2 (8.25 + 2.75)
22= 2 (11)
22= 22
Therefore, having Width = 2.75 is correct because it will arrived to a Perimeter of 22 as shown in th equation. Good Job!